Can you please solve the below algebra problem and explain the steps for me. Please and Thankyou.
if y=sqrt (x-2) fill in the following table for x=0,1,2,3,4. round to three decimal places where necessary.

Answer by dfrazzetto(283)


y= sqrt(x-2)
y=sqrt(0-2) = sqrt(-2) = 1.414i (imaginary)
y=sqrt(1-2) = sqrt(-1) = i (imaginary)
y=sqrt(2-2) = sqrt(0) = 0
y=sqrt(3-2) = sqrt(1) = 1
y=sqrt(4-2) = sqrt(2) = 1.414
Steps: used square root button on calculator where needed